# A generalization of the Romanoff theorem

### v2

@@ 1. Introoduction
Erdős paid much attention Romanoff’s theorem and the proof of it was later simplified by Erdős and Turán [11] and Erdős [9]. In [10], Erdős extended powers of $2$ in the Romanoff theorem to a more general sequence $\{a_{k}\}$. Precisely, let $a_{1}<a_{2}<\cdots$ be an infinite sequence of positive integers satisfying $a_{k}|a_{k+1}$. Then Erdős proved that the necessary and sufficient condition that the sequence $p+a_{k}$ should have positive density is that
@@
As an application, Yang and Chen obtained the following estimates
@@
In a prior article, the author [6] completely solved Problem 2 by showing that $k=2$ is admissible. Thus,
@@
In a subsequent article, the author [7] answered Problem 1 negatively and proved that for any subset $\mathcal{B}$ of $c$-condition with $\mathcal{B}(x)\gg x^{\alpha}$ we have
@@
I also pointed out that the above lower bound is sharp via concrete examples. Following the line of Problem 2, Chen and Xu [3] obtained Theorem 1 with additional condition: ‘at least one of $r_{1},...,r_{t}$ is an integer’. And in [3, Remark 1], they commented that ‘We believe that it is unnecessary that at least one of $r_{1},...,r_{t}$ is an integer in our theorem’. It is just our purpose to remove this additional condition which forms a clean generalization (Theorem 1) of the Romanoff theorem.
@@
The main difference of the proof is that when we encounter Eq. (9), we can not remove the floor function directly since $r_{1}$ will not be an integer any more (comparing this with the corresponding argument of Chen and Xu [3, Theorem 3.1]). By an elementary handle we transform Eq. (9) to Eq. (10) and then a result of Erdős and Turán could be used therein. This finite form result on uniform distribution is crucial here since the modulo $e_{2}(d)$ could be significantly greater than the range of $k_{1}$ (see Eq. (9)), which is different to the usual researches on Piatetski-Shapiro like sequences.
Throughout the proof, $\mathcal{S}(x)$ denotes the number of elements of $\mathcal{S}$ not exceeding $x$.
@@
$A(x)\mathcal{P}(x)\ll x$
via Eq. (2) and the prime number theorem. Thus, to complete our proof, it would suffice to prove that
@@
For a given $g\leq x$, suppose first that there is at most one solution $k_{1}$ to the congruence equation
$2^{\left\lfloor k_{1}^{r_{1}}\right\rfloor}\equiv g\ ({\rm{mod}}\ d),$ (7)
then for those $d$ the double sums in right–hand side of Eq. (2) are bounded by
$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\leq\prod_{2<p<\log x}\left(1+\frac{1}{p}\right)\ll\log\log x$ (8)
which is admissible for our purpose, where the last estimate follows from the famous theorem of Mertens. We are left over to consider the remaining $d$ for which there are at least two solutions $k_{1}$ satisfying Eq. (7). In this case, let $k_{1,0}$ be such a solution. Then we have
@@
for some $10\leq\ell\leq e_{2}(d)-$ 1, where $e_{2}(d)$ is the least positive integer $m$ so that $d|2^{m}-1$. It follows that
$e_{2}(d)f+\ell\leq k_{1}^{r_{1}}<e_{2}(d)f+\ell+1$
for some $f\in\mathbb{Z}$ from Eq. (9) and vice versa, which clearly means that
$\frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}<\frac{\ell+1}{e_{2}(d)},$
where $\{y\}$ stands for the fractional part of $y$. It would be more pleasant to relax the above inequality to
$\frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}\leq\frac{\ell+1}{e_{2}(d)}$ (10)
and we shall consider Eq. (10) from now on. We are in a position to employ a ‘finite’ form regarding the classical theorem of Weyl [24] on uniform distribution which was later established by Erdős and Turán [12, Theorem III].
Lemma 1. If $\varphi_{1},\varphi_{2},\cdot\cdot\cdot,\varphi_{n}$ are real satisfying that
$\bigg{|}\sum_{v=1}^{n}e^{ki\varphi_{v}}\bigg{|}\leq\psi(k)\quad(k=1,2,...,m,...),$
then for $m\geq 1$ and arbitrary $0\leq\alpha<\beta\leq 2\pi$ we have
[…]
with a numerical constant $C$.
Now we take $\varphi_{v}=2\pi\left\{\frac{v^{r_{1}}}{e_{2}(d)}\right\},n=\left\lfloor\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\right\rfloor,\alpha=\frac{2\pi\ell}{e_{2}(d)}$ and $\beta=\frac{2\pi(\ell+1)}{e_{2}(d)}$ in Lemma 1, then there is some absolute constant $C$ so that
$\Bigg{|}\sum_{\begin{subarray}{c}k_{1}\leq\left(\log 2x/\log 2\right)^{1/r_{1}}\\ \frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}\leq\frac{(\ell+1)}{e_{2}(d)}\end{subarray}}\!\!\!\!1-\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{e_{2}(d)}\Bigg{|}<C\left(\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{m+1}+\sum_{k=1}^{m}\frac{\psi(k)}{k}\right),$ (11)
where $m\geq 1$ is a parameter to be decided later and $\psi(k)$ is the quantity satisfying
$\bigg{|}\sum_{v=1}^{\left(\log 2x/\log 2\right)^{1/r_{1}}}e^{ki2\pi\left\{\frac{v^{r_{1}}}{e_{2}(d)}\right\}}\bigg{|}\leq\psi(k).$ (12)
To give an appropriate estimate of $\psi(k)$, we use the van der Corput type bounds [23, Theorems 5.9, 5.11 and 5.13].
Lemma 2. Let $f(x)$ be real and have continuous derivates up to the $j$–th order, where $j\geq 2$. Let $\lambda_{j}\leq f^{(j)}(x)\leq h\lambda_{j}$ (or the same for $-f^{(j)}(x)$). Let $b-a\geq 1,J=2^{j-1}$. Then
$\sum_{a<n\leq b}e^{2\pi if(n)}=O\left(h^{2/J}(b-a)\lambda_{j}^{1/(2J-2)}\right)+O\left((b-a)^{1-2/J}\lambda_{j}^{-1/(2J-2)}\right),$
where the constants implied by big-$O$ are independent of $j$.
Let $f(x)=\frac{kx^{r_{1}}}{e_{2}(d)}$ and $r$ be the least integer greater than $r_{1}$. Then
$f^{(r)}(x)=\frac{k}{e_{2}(d)}r_{1}(r_{1}-1)\cdot\cdot\cdot(r_{1}-r+1)x^{r_{1}-r}.$
Considering with $a\leq x<b\leq 2a$, we clearly have
$\frac{kC_{1}}{e_{2}(d)(2a)^{r-r_{1}}}\leq f^{(r)}(x)\leq 2^{r-r_{1}}\frac{kC_{1}}{e_{2}(d)(2a)^{r-r_{1}}},$
where $C_{1}=r_{1}(r_{1}-1)\cdot\cdot\cdot(r_{1}-r+1)$. Employing Lemma 2 we obtain that
$\sum_{a<n\leq b}e^{2\pi i\frac{kn^{r_{1}}}{e_{2}(d)}}$ $\ll_{r_{1}}a\left(\frac{k}{e_{2}(d)a^{r-r_{1}}}\right)^{1/(2R-2)}+a^{1-2/R}\left(\frac{k}{e_{2}(d)a^{r-r_{1}}}\right)^{-1/(2R-2)}$
$\ll a^{1-\frac{r-r_{1}}{2R-2}}\frac{k^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+a^{1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}}\frac{e_{2}(d)^{\frac{1}{2R-2}}}{k^{\frac{1}{2R-2}}},$
where $R=2^{r-1}$. The dichotomy of $a$ between the interval $\left[1,\left(\log 2x/\log 2\right)^{1/r_{1}}\right]$ yields
$\sum_{v=1}^{\left(\log 2x/\log 2\right)^{1/r_{1}}}e^{ki2\pi\left\{\frac{v^{r_{1}}}{e_{2}(d)}\right\}}$ $=\sum_{v=1}^{\left(\log 2x/\log 2\right)^{1/r_{1}}}e^{ki2\pi\frac{v^{r_{1}}}{e_{2}(d)}}$
$\ll_{r_{1}}\left(\log x\right)^{\frac{1}{r_{1}}\left(1-\frac{r-r_{1}}{2R-2}\right)}\frac{k^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+(\log x)^{\frac{1}{r_{1}}\left(1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}\right)}\frac{e_{2}(d)^{\frac{1}{2R-2}}}{k^{\frac{1}{2R-2}}}.$
Thus, by Eq. (12) we can choose $\psi(k)$ to satisfy that
$\psi(k)\ll_{r_{1}}\left(\log x\right)^{\frac{1}{r_{1}}\left(1-\frac{r-r_{1}}{2R-2}\right)}\frac{k^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+(\log x)^{\frac{1}{r_{1}}\left(1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}\right)}\frac{e_{2}(d)^{\frac{1}{2R-2}}}{k^{\frac{1}{2R-2}}},$
from which it follows directly via partial summations that
$\sum_{k=1}^{m}\frac{\psi(k)}{k}\ll_{r_{1}}\left(\log x\right)^{\frac{1}{r_{1}}\left(1-\frac{r-r_{1}}{2R-2}\right)}\frac{m^{\frac{1}{2R-2}}}{e_{2}(d)^{\frac{1}{2R-2}}}+(\log x)^{\frac{1}{r_{1}}\left(1-\frac{2}{R}+\frac{r-r_{1}}{2R-2}\right)}e_{2}(d)^{\frac{1}{2R-2}}.$ (13)
We take $m=\lfloor\log\log x\rfloor$ and then insert Eq. (13) into Eq. (11). Then
$\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ \frac{\ell}{e_{2}(d)}\leq\left\{\frac{k_{1}^{r_{1}}}{e_{2}(d)}\right\}\leq\frac{(\ell+1)}{e_{2}(d)}\end{subarray}}\!\!\!\!1=\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{e_{2}(d)}+O_{r_{1}}\left(\frac{\left(\log 2x/\log 2\right)^{1/r_{1}}}{\log\log x}+\sum_{k=1}^{\lfloor\log\log x\rfloor}\frac{\psi(k)}{k}\right).$ (14)
@@
Collecting Eqs. (2), (8), (13) and (14), we conclude from the analysis above that
@@
$\sum_{\begin{subarray}{c}a_{1},a_{2}\in\mathcal{A}\\ a_{1}<a_{2}\leq x\end{subarray}}\sum_{\begin{subarray}{c}d|a_{2}-a_{1}\\ 2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\!\!\frac{\mu^{2}(d)}{d}\ll_{r_{1}}(\log x)^{2-\frac{1}{r_{1}}}\log\log x+S_{1}(x)+S_{2}(x)+S_{3}(x)+S_{4}(x)$ (15)
@@
$S_{1}(x):=(\log x)^{2}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{de_{2}(d)},\qquad S_{2}(x):=\frac{(\log x)^{2}}{\log\log x}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d},$
$S_{3}(x):=(\log x)^{2-\frac{r-r_{1}}{r_{1}(2R-2)}}(\log\log x)^{{}^{\frac{1}{2R-2}}}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{de_{2}(d)^{\frac{1}{2R-2}}}$
@@
$S_{4}(x):=(\log x)^{2-\left(\frac{2}{R}-\frac{r-r_{1}}{r_{1}(2R-2)}\right)}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)e_{2}(d)^{\frac{1}{2R-2}}}{d}.$
@@
for any $\varepsilon>0$. Thus, we deduce that $S_{1}(x)\ll(\log x)^{2}$ and
$S_{3}(x)\ll(\log x)^{2-\frac{r-r_{1}}{r_{1}(2R-2)}}(\log\log x)^{{}^{\frac{1}{2R-2}}}\ll(\log x)^{2}.$
By Eq. (8), we also have $S_{2}(x)\ll(\log x)^{2}$. For $2\nmid d$, we have
$2^{\varphi(d)}\equiv 1\ ({\rm{mod}}\ d)$
by the Euler theorem, where $\varphi(\cdot)$ is the Euler totient function. So $e_{2}(d)\leq\varphi(d)\leq d$, from which it follows that
$S_{4}(x)\leq(\log x)^{2-\left(\frac{2}{R}-\frac{r-r_{1}}{r_{1}(2R-2)}\right)}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d^{1-\frac{1}{2R-2}}}.$
Changing back the sum to product, we get
$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d^{1-\frac{1}{2R-2}}}\leq\prod_{2<p<\log x}\left(1+\frac{1}{p^{1-\frac{1}{2R-2}}}\right)=\exp\left(\sum_{2<p<\log x}\log\left(1+\frac{1}{p^{1-\frac{1}{2R-2}}}\right)\right).$
Employing the elementary inequality $\log(1+z)<z$ for any $z>0$, the right–hand side above is
$<\exp\left(\sum_{2<p<\log x}\frac{1}{p^{1-\frac{1}{2R-2}}}\right).$
Using the prime number theorem as well as the partial summations, we have
$\sum_{2<p<\log x}\frac{1}{p^{1-\frac{1}{2R-2}}}\ll\frac{(\log x)^{\frac{1}{2R-2}}}{\log\log x},$
which implies $S_{4}(x)\ll(\log x)^{2}$ on noting that
$\frac{2}{R}-\frac{r-r_{1}}{r_{1}(2R-2)}>\frac{1}{2R-2}.$
Therefore, the right–hand side of Eq. (15) is bounded by $(\log x)^{2}$ which proves Eq. (5) and hence Eq. (3). ∎

### v3

@@ 1. Introoduction
Erdős paid much attention to Romanoff’s theorem and the proof of it was later simplified by Erdős and Turán [11] and Erdős [9]. In [10], Erdős extended powers of $2$ in the Romanoff theorem to a more general sequence $\{a_{k}\}$. Precisely, let $a_{1}<a_{2}<\cdots$ be an infinite sequence of positive integers satisfying $a_{k}|a_{k+1}$. Then Erdős proved that the necessary and sufficient condition that the sequence $p+a_{k}$ should have positive density is that
@@
As an application, Yang and Chen obtained the following estimate
@@
In a prior article, the first named author [6] completely solved Problem 2 by showing that $k=2$ is admissible. Thus,
@@
In a subsequent article, the first named author [7] answered Problem 1 negatively and proved that for any subset $\mathcal{B}$ of $c$-condition with $\mathcal{B}(x)\gg x^{\alpha}$ we have
@@
The first named author also pointed out that the above lower bound is sharp via concrete examples. Following the line of Problem 2, Chen and Xu [3] obtained Theorem 1 with additional condition: ‘at least one of $r_{1},...,r_{t}$ is an integer’. And in [3, Remark 1], they commented that ‘We believe that it is unnecessary that at least one of $r_{1},...,r_{t}$ is an integer in our theorem’. It is just our purpose to remove this additional condition which forms a clean generalization (Theorem 1) of the Romanoff theorem.
@@ 2. Proofs
Throughout the proof, $\mathcal{S}(x)$ denotes the number of elements of $\mathcal{S}$ not exceeding $x$ and $x$ is supposed to be sufficiently large.
@@
$\mathcal{A}(x)\mathcal{P}(x)\ll x$
via Eq. (2) and the prime number theorem. Thus, to complete our proof, it suffices to prove that
@@
For some integers $d$, suppose first that the number of $k_{1}$ satisfying the congruence
$2^{\left\lfloor k_{1}^{r_{1}}\right\rfloor}\equiv g\ ({\rm{mod}}\ d)$ (7)
is not large than $(\log x)^{1/r_{1}}/\log\log x$, then it is admissible for those $d$ since the double sums in right–hand side of Eq. (2) are bounded now by
$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\frac{(\log x)^{1/r_{1}}}{\log\log x}$ $\leq\frac{(\log x)^{1/r_{1}}}{\log\log x}\prod_{2<p<\log x}\left(1+\frac{1}{p}\right)$
$=\frac{(\log x)^{1/r_{1}}}{\log\log x}\exp\left(\sum_{2<p<\log x}\log\left(1+\frac{1}{p}\right)\right)$
$<\frac{(\log x)^{1/r_{1}}}{\log\log x}\exp\left(\sum_{2<p<\log x}\frac{1}{p}\right)$
$\ll(\log x)^{1/r_{1}},$ (8)
where the last estimate follows from the famous theorem of Mertens. We are left over to consider the remaining $d$ for which the number of $k_{1}$ satisfying Eq. (7) is greater than $(\log x)^{1/r_{1}}/\log\log x$. In this case, let $k_{1,0}$ be such a solution. Then we have
@@
for some $10\leq\ell\leq e_{2}(d)-$ 1, where $e_{2}(d)$ is the least positive integer $m$ so that $d|2^{m}-1$. Furthermore, in case that (9) has more than $(\log x)^{1/r_{1}}/\log\log x$ solutions of $k_{1}$, we claim that
$e_{2}(d)\ll_{r_{1}}(\log x)^{1-\frac{1}{r_{1}}}\log\log x,$ (10)
which would be one of the key points in the following arguments. In fact, suppose that $L$ is the minimal distance between two consecutive solutions $k_{1}^{\prime}$ and $k_{1}^{\prime}+L$ of the congruence equation (9) with $k_{1}^{\prime}>(\log x)^{\frac{1}{2r_{1}}}$, then the number of solutions $k_{1}$ to (9) is not more than
$(\log x)^{\frac{1}{2r_{1}}}+\frac{(\log x)^{1/r_{1}}}{L}+1.$
Hence, by our assumption on the number of the solutions, we have
$(\log x)^{\frac{1}{2r_{1}}}+\frac{(\log x)^{1/r_{1}}}{L}+1\geq\frac{(\log x)^{1/r_{1}}}{\log\log x},$
which clearly means that $L\leq 2\log\log x.$ By the definition of $L$, we also have
$\lfloor(k_{1}^{\prime}+L)^{r_{1}}\rfloor-\lfloor(k_{1}^{\prime})^{r_{1}}\rfloor\geq e_{2}(d),$
from which it follows that
$e_{2}(d)\leq(k_{1}^{\prime}+L)^{r_{1}}-(k_{1}^{\prime})^{r_{1}}+1\ll_{r_{1}}(k_{1}^{\prime})^{r_{1}-1}L\ll(\log x)^{1-1/r_{1}}\log\log x.$
@@
We now conclude from (2), (9), (10) and the above proofs that we only need to show
@@
$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ \lfloor k_{1}^{r_{1}}\rfloor\equiv\ell\ ({\rm{mod}}\ e_{2}(d))\end{subarray}}1\ll(\log x)^{\frac{1}{r_{1}}}.$ (11)
For any $k_{1}\leq(\log x)^{1/r_{1}}$ satisfies (9), we clearly have
$e_{2}(d)q+\ell\leq k_{1}^{r_{1}}<e_{2}(d)q+\ell+1$
for some nonnegative integer $q$, which is equivalent to
$\frac{k_{1}^{r_{1}}-\ell-1}{e_{2}(d)}<q\leq\frac{k_{1}^{r_{1}}-\ell}{e_{2}(d)}.$ (12)
It is plain that an integer $q$ satisfies (12) if and only if
$\left\lfloor\frac{k_{1}^{r_{1}}-\ell}{e_{2}(d)}\right\rfloor-\left\lfloor\frac{k_{1}^{r_{1}}-\ell-1}{e_{2}(d)}\right\rfloor=1,$
from which it follows that
$\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ \lfloor k_{1}^{r_{1}}\rfloor\equiv\ell\ ({\rm{mod}}\ e_{2}(d))\end{subarray}}1$ $\leq\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\end{subarray}}\bigg{(}\left\lfloor\frac{k_{1}^{r_{1}}-\ell}{e_{2}(d)}\right\rfloor-\left\lfloor\frac{k_{1}^{r_{1}}-\ell-1}{e_{2}(d)}\right\rfloor\bigg{)}$
$=\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\end{subarray}}\bigg{(}\frac{1}{e_{2}(d)}-\psi\left(\frac{k_{1}^{r_{1}}-\ell}{e_{2}(d)}\right)-\psi\left(\frac{k_{1}^{r_{1}}-\ell-1}{e_{2}(d)}\right)\bigg{)}$
$\ll_{r_{1}}\frac{(\log x)^{1/r_{1}}}{e_{2}(d)}+O\bigg{(}\max_{|\theta|\leq 1}\bigg{|}\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\end{subarray}}\psi\left(\frac{k_{1}^{r_{1}}}{e_{2}(d)}+\theta\right)\bigg{|}\bigg{)},$ (13)
where $\psi(t)=\{t\}-1/2$ is the sawtooth function. Now, we write
$\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\end{subarray}}\psi\left(\frac{k_{1}^{r_{1}}}{e_{2}(d)}+\theta\right)=S_{1}+S_{2},$ (14)
@@
$S_{1}=\sum_{k\leq e_{2}(d)^{1/r_{1}}}\psi\left(\frac{k_{1}^{r_{1}}}{e_{2}(d)}+\theta\right)\quad\text{and}\quad S_{2}=\sum_{e_{2}(d)^{1/r_{1}}<k\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}}\psi\left(\frac{k_{1}^{r_{1}}}{e_{2}(d)}+\theta\right).$
Trivially, we have
$S_{1}\ll e_{2}(d)^{1/r_{1}}.$ (15)
We are now in a position to use the following van der Corput type bounds [14, Lemma 4.3] for a suitable estimate of $S_{2}$.
Lemma 1. For any real number $t,$ let $\psi(t)=\{t\}-1/2$. Suppose $K\geq 3$ is a real number, $Y>0.$ Suppose $\alpha>0~{}(\alpha\notin\mathbb{Z})$ is a fixed real number. Then for any exponent pair $(\kappa,\lambda)$ we have
$\sum_{K\leq k<2K}\psi\left(Yk^{\alpha}+\theta\right)\ll\frac{K^{1-\alpha}}{Y}+Y^{\frac{\kappa}{1+\kappa}}K^{\frac{\lambda+\kappa\alpha}{1+\kappa}}.$
For $e_{2}(d)^{1/r_{1}}\leq K\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}$, by Lemma 1 with $Y=\frac{1}{e_{2}(d)}$ we have
$\sum_{K\leq k_{1}<2K}\psi\left(\frac{k_{1}^{r_{1}}}{e_{2}(d)}+\theta\right)$ $=\sum_{K\leq k<2K}\psi\left(Yk_{1}^{r_{1}}\right)$
$\ll\frac{1}{YK^{r_{1}-1}}+Y^{\frac{\kappa}{1+\kappa}}K^{\frac{r_{1}\kappa+\lambda}{1+\kappa}}$
$\ll\frac{e_{2}(d)}{K^{r_{1}-1}}+e_{2}(d)^{-\frac{\kappa}{1+\kappa}}K^{\frac{r_{1}\kappa+\lambda}{1+\kappa}}$
since $Y=1/e_{2}(d)$, which combining a splitting argument yields
$S_{2}=\sum_{e_{2}(d)^{1/r_{1}}<k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}}\psi\left(\frac{k^{r_{1}}}{e_{2}(d)}+\theta\right)\ll(\log x)^{\frac{1}{r_{1}}\cdot\frac{r_{1}\kappa+\lambda}{1+\kappa}}e_{2}(d)^{-\frac{\kappa}{1+\kappa}}+e_{2}(d)^{1/r_{1}}.$ (16)
Taking (15) and (16) into (14), we have
$\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ \lfloor k_{1}^{r_{1}}\rfloor\equiv\ell\ ({\rm{mod}}\ e_{2}(d))\end{subarray}}1\ll\frac{(\log x)^{1/r_{1}}}{e_{2}(d)}+(\log x)^{\frac{1}{r_{1}}\cdot\frac{r_{1}\kappa+\lambda}{1+\kappa}}e_{2}(d)^{-\frac{\kappa}{1+\kappa}}+e_{2}(d)^{1/r_{1}}.$
Therefore, we have
$\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\sum_{\begin{subarray}{c}k_{1}\leq\left(\frac{\log 2x}{\log 2}\right)^{1/r_{1}}\\ \lfloor k_{1}^{r_{1}}\rfloor\equiv\ell\ ({\rm{mod}}\ e_{2}(d))\end{subarray}}1\ll_{r_{1}}\mathcal{W}_{1}+\mathcal{W}_{2}+\mathcal{W}_{3},$ (17)
where
$\mathcal{W}_{1}=\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}\frac{(\log x)^{1/r_{1}}}{e_{2}(d)},$
$\mathcal{W}_{2}=\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}(\log x)^{\frac{1}{r_{1}}\cdot\frac{r_{1}\kappa+\lambda}{1+\kappa}}e_{2}(d)^{-\frac{\kappa}{1+\kappa}}$
@@
$\mathcal{W}_{3}=\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}e_{2}(d)^{1/r_{1}}.$
@@
for any $\varepsilon>0$. Using (18), we know that
$\mathcal{W}_{1}\leq(\log x)^{1/r_{1}}\sum_{\begin{subarray}{c}2\nmid d\end{subarray}}\frac{1}{de_{2}(d)}\ll(\log x)^{1/r_{1}}$
and together with the estimate given in (2), we get
$\mathcal{W}_{3}$ $\ll\Big{(}(\log x)^{1-1/r_{1}}\log\log x\Big{)}^{1/r_{1}}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\\ e_{2}(d)\leq(\log x)^{1-1/r_{1}}\log\log x\end{subarray}}\frac{\mu^{2}(d)}{d}$
$\ll(\log x)^{\frac{1}{r_{1}}-\frac{1}{r_{1}^{2}}}(\log\log x)^{\frac{1}{r_{1}}}\sum_{\begin{subarray}{c}d\leq x,~{}2\nmid d\\ P^{+}(d)<\log x\end{subarray}}\frac{\mu^{2}(d)}{d}$
$\ll(\log x)^{\frac{1}{r_{1}}-\frac{1}{r_{1}^{2}}}(\log\log x)^{\frac{1}{r_{1}}+1}$
$\ll(\log x)^{\frac{1}{r_{1}}}.$
On noting (17), to establish the estimates of (11) it remains to show $\mathcal{W}_{2}\ll(\log x)^{\frac{1}{r_{1}}}$. Again, by (18) and the estimates in (2), it suffices to choose a suitable exponent pair $(\kappa,\lambda)$ such that
$\frac{r_{1}\kappa+\lambda}{1+\kappa}<1.$ (19)
For any integer $q\geq 1$, we know that
$\left(\frac{1}{4Q-2},1-\frac{q+1}{4Q-2}\right)$
is an exponent pair, thanks to the display of [14, Page 60], where $Q=2^{q}$. We now take $q=\lfloor r_{1}\rfloor+1$ and then
$r_{1}\kappa+\lambda=\frac{r_{1}}{4Q-2}+1-\frac{q+1}{4Q-2}<1,$
[…]